# A Fast, Analytic Empirical Model of the Gaia Data Release 3 Astrometric Orbit Catalog Selection Function

### v1

@@ 6.4 Gaia DR4
As of late 2024, Gaia DR4 is projected to be released in mid-2026. DR4 will nearly double the time baseline of DR3 (66 months of data vs. 34 months). It is expected that the parallax and photocenter orbit semimajor axis uncertainties will be reduced by a factor of $\sqrt{2}$ due to having double the number of observations, and proper motion errors will be reduced by a factor of $\sqrt{8}$, with a factor of 2 in the doubled baseline, and another factor of $\sqrt{2}$ for the doubled number of observations.
It is beyond the scope of this work to extend this forward model framework to forecast for Gaia DR4. Several parts of this forward model, such as the empirical removal of solutions at certain orbital periods (§4.3), would require knowledge of the DR4 processing pipeline to forecast. An alternative way to create a conservative forecast of the selection of DR4 binaries is to model per-epoch photometry and assume that the processing cascade remains the same as that of DR3 (El-Badry et al., 2024). However, once DR4 is released, it would not be difficult to extend the model presented here.
@@
6.5 Comparison to El-Badry et al. (2024)
In order to characterize the selection function of the Gaia DR3 NSS astrometric orbits catalog, El-Badry et al. (2024) forward model individual epoch observations and validate it using a binary population synthesis model. Many of the underlying components of the population synthesis are similar or identical to that described in this paper; they also use galaxia and COSMIC to simulate the population of binary stars within 2 kpc. Minor differences include the extinction ($A_{G}=2.3E(B-V)$ in this work vs. $A_{G}=2.8E(B-V)$ in El-Badry et al. 2024), number of binaries at birth within 2 kpc (170 million in this work vs. 195 million in El-Badry et al. 2024), and the white dwarf initial-final mass relation (Kalirai et al. 2008 in this work vs. Weidemann 2000 in El-Badry et al. 2024).
To generate the individual observations, they use GOST to obtain the scan times and scan angles, and use Holl et al. (2023) for the noise model of the observations. They then fit these mock observations following the cascade described in Halbwachs et al. (2023). The advantage of modeling the observations at the epoch level is that each step described in the processing pipeline can be modeled. The main drawback is the computational expense. It requires approximately $10-20$ seconds to run the full astrometric cascade of El-Badry et al. (2024) for each source that receives an orbital solution, or about 5000 CPU hours to produce a full mock DR3 binary catalog. In contrast, it requires approximately 2 seconds to run 100,000 evaluations of the empirical forward model of this work at a fixed position on sky; most of that time comes from querying the 3-D dust map.
We also compare our results for the inclusion probability of Gaia BH1, BH2, and BH3 in the Gaia NSS astrometric orbits catalog. The results agree well for Gaia BH1 (33% in this work, and 39% in El-Badry et al. (2024)). For Gaia BH2 and BH3, El-Badry et al. (2024) find significantly higher inclusion probabilities (48% and 10%, as compared to the 4% and 0% in this work). One difference is that the values for El-Badry et al. are sky-averaged, while the values presented here are a function of (RA, Dec). For Gaia BH1, the detection probability would be 2×2$\times$2 × higher at higher ecliptic latitudes (Figure 9); for Gaia BH2, the detection probability does not vary across the sky (Figure 9). El-Badry et al. (2024) also calculate the change in inclusion probabilities as a function of distance. For Gaia BH1, the results are qualitatively similar, although our inclusion probability rises more steeply toward closer distances. For Gaia BH2, the results disagree, mainly because our inclusion probability does not exceed $\approx 4\%$. This may be due to strong covariances between $P_{orb}$, the mass of the dark companion, and detection probability, which is not captured in the probabilistic model that removes long orbital period solutions.
@@ 7 Conclusions
To address this, we create a framework to forward model the selection function of the Gaia DR3 NSS astrometric orbits catalog. Using a binary population synthesis model, we validate this framework. We find the method produces reasonable agreement with the observed distribution of periods, parallaxes, photocenter sizes, ruwe, magnitudes, and parallax errors; the one parameter that is substantially discrepant is the eccentricity; we suspect the assumed birth population of eccentricities is the culprit.
As an example application of this framework, we use it the model to calculate the probability of Gaia BH1, BH2, and BH3 being included in the NSS astrometric orbits catalog. We find that Gaia BH1 has a relatively high chance of being included, but Gaia BH2 was a much luckier detection. Gaia BH3 was unlikely to be detected, due to its very long orbital period.
@@
Appendix A Parallax and photocenter orbit semimajor axis error model
Here we investigate the relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$, and $\sigma_{AL}/\sqrt{N_{vis}}$. There is a linear relationship between $\sigma_{\varpi}$, $\sigma_{a_{0}}$ and $\sigma_{AL}/\sqrt{N_{vis}}$. Figure 12 shows this relationship, binned by the magnitude of the source and the number of visibility periods used. The binned value is indicated by the dot, the error bar is the error on the mean of that bin. Each dot represents a visibility period for that magnitude range; they are connected in order of decreasing number of visibility periods. The relationship gets noisy towards fainter magnitudes ($G\gtrsim 16$) and hence the line connecting the dot is no longer straight. The relationship for $138<G<$ 13 are all similar because $\sigma_{AL}$ is roughly constant in this magnitude range. However, it is unclear why going to fainter magnitudes, $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ are smaller for a given $\sigma_{AL}/\sqrt{N_{vis}}$. To investigate, we make the corresponding plot in $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ but using a random subset of the gaia source catalog. We find a similar relation persists, albeit the discrepancy at fixed $\sigma_{AL}/\sqrt{N_{vis}}$ is not as pronounced. We thus suspect that there is a magnitude dependence on $\sigma_{\varpi}$ and $\sigma_{a_{0}}$ that is not necessarily entirely captured in $\sigma_{AL}$.
Figure 12: $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (top panel), $\sigma_{a_{0}}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for the NSS astrometric orbit sources (middle panel), and $\sigma_{\varpi}$ vs. $\sigma_{AL}/\sqrt{N_{vis}}$ for a random subset of gaia_source sources (bottom panel). The sources are binned by magnitude and number of visibility periods used; for a given magnitude, the points are connected by a line indicated increasing number of visibility periods. The NSS catalogs exhibit a discontinuity between different magnitudes not seen in the gaia_source catalog.
Appendix B Cross-match to external catalogs

### v2

@@ 6.4 Gaia DR4
As of 2025, Gaia DR4 is projected to be released in late 2026. DR4 will nearly double the time baseline of DR3 (66 months of data vs. 34 months). It is expected that the parallax and photocenter orbit semimajor axis uncertainties will be reduced by a factor of $\sqrt{2}$ due to having double the number of observations, and proper motion errors will be reduced by a factor of $\sqrt{8}$, with a factor of 2 in the doubled baseline, and another factor of $\sqrt{2}$ for the doubled number of observations.
It is beyond the scope of this work to extend this forward model framework to forecast for Gaia DR4. In principle, this can by done by re-generating the catalogs of El-Badry et al. (2024) using the 66 month DR4 baseline, and re-calibrating the empirical corrections here.
@@ 7 Conclusions
To address this, we create a fast, simple, and physically interpretable framework to forward model the selection function of the Gaia DR3 NSS astrometric orbits catalog. Our method combines analytic prescriptions with detailed results from El-Badry et al. (2024). Using a binary population synthesis model, we validate this framework. We find the method produces reasonable agreement with the observed distribution of periods, parallaxes, photocenter sizes, ruwe, magnitudes, and parallax errors; the one parameter that is substantially discrepant is the eccentricity; we suspect the assumed birth population of eccentricities is the culprit. We find the proposed Rayleigh distribution of Wu et al. (2025) to better explain the eccentricity distribution of Gaia sources.
As an example application of this framework, we use it the model to calculate the probability of Gaia BH1 and BH2 being included in the NSS astrometric orbits catalog. We find that both Gaia BH1 and BH2 had relatively high probabilities ($>20$%) of being included in the Gaia DR3 NSS astrometric orbits catalog, and that there are a few thousand Sun-like star + BH binaries in the entire Galaxy.
@@
Appendix A Eccentricity error $\sigma_{e}$
Table 1 of Halbwachs et al. (2023) lists the criteria for acceptance as an orbital solution as $F_{2}<25$, $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$, $a_{0}/\sigma_{a_{0}}>\max(5,158/\sqrt{P_{orb}/\textrm{days}})$, and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$. As described in §5.3 “Selection of the orbital solutions” of Halbwachs et al. (2023), the criteria $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ only remove an additional 7% of solutions after the other three criteria were applied. Here, we provide justification to ignore $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ in our forward modeling.
We analyze all the orbital solutions in the simulations of El-Badry et al. (2024) that pass the filters $\varpi/\sigma_{\varpi}>20,000\textrm{ day}/P_{orb}$ and $a_{0}/\sigma_{a_{0}}>5$. We then 1) remove solutions that also pass the filter $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$, 2) remove solutions that pass the $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filter, and 3) remove solutions that pass both $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filters. We show the distribution of $P_{orb}$, $e$, and $a_{0}$ for these solutions in Figure 14. The effect of excluding the $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ filter makes almost no difference on the inferred parameter distributions.
Figure 14: Distributions of the orbital period $P_{orb}$, eccentricity $e$, and photocenter orbit semimajor axis $a_{0}$ depending on whether the filters on $a_{0}/\sigma_{a_{0}}>158/\sqrt{P_{orb}/\textrm{days}}$ and $\sigma_{e}<0.079\ln{P_{orb}/\textrm{d}}-0.244$ were included.
Appendix B Goodness-of-fit $F_{2}$
The $F_{2}>25$ criterion is mainly for removing spurious solutions. Because our modeling assumes the true orbital parameters are known and that all systems are binaries (and not, e.g., triple systems), these spurious solutions do not exist in our forward model. We validate this using the simulations in El-Badry et al. (2024). Figure 15 shows period vs. eccentricity for sources where an orbital solution is tried, with bad solutions ($F_{2}>25$) in orange points. The left panel shows the true period and eccentricity (which are known because they are simulation inputs). The right panel are the inferred values performed from fitting simulated Gaia data (c.f. the middle panel of Figure 9 of El-Badry et al. 2024). The solutions removed by the $F_{2}>25$ criterion are all artifacts due to incorrect fits and have true orbital periods $P_{orb}>10,000$ days. Given that we only model sources with $P_{orb}<1500$ days, we are justified in assuming that $F_{2}<25$ is always satisfied for our sources.
Figure 15: Spurious solutions as identified in the epoch modeling of El-Badry et al. (2024). Sources with $F_{2}>25$ (orange) tend to have incorrectly measured periods. The inferred periods (right panel) are much shorter than the true periods ($P_{orb}\gtrsim 10,000$ days), which Gaia DR3 is insensitive to.
Appendix C Acceleration modeling and orbital properties
Here we show the correlations between the criteria for accepting an acceleration model and their dependence on the true orbital properties. Figure 16 demonstrates the relationship between the significances of the 7- and 9-parameter acceleration solutions $s_{7},s_{9}$ and the parallax over errors $\varpi/\sigma_{\varpi}$ of the respective models. As described in §2.3.2 and §4.3 of Halbwachs et al. (2023), a 9-parameter solution must have $\varpi/\sigma_{\varpi}>2.1s_{9}^{1.05}$, $F_{2}<25$, and $s_{9}>12$ to be provisionally accepted in the main processing; a 7-parameter solution must have $\varpi/\sigma_{\varpi}>1.2s_{7}^{1.05}$, $F_{2}<25$, and $s_{7}>12$. Note that $\varpi/\sigma_{\varpi}$ is different for each different model (7-parameter acceleration, 9-parameter acceleration, orbital model), as different values are inferred when being fit with different NSS models.
Figure 17 shows the probability of a source to be accepted into the 9-parameter and 7-parameter acceleration model branch of the cascade, as a function of the orbital period and photocenter orbit semimajor axis size. The left panel of Figure 17 shows the probability that a source with $\texttt{ruwe}>1.4$ is a 9-parameter acceleration solution candidate as a function of $P_{orb}$ and $a_{0}/\sigma_{AL}$. The right panel of Figure 17 shows the probability that a source that has $\texttt{ruwe}>1.4$ and that is not a candidate for a 9-parameter acceleration solution is a candidate for a 7-parameter acceleration solution. In both cases, longer period solutions $P_{orb}\gtrsim 500$ days are more likely to have an accepted acceleration solution. For 9-parameter acceleration solutions, intermediate $a_{0}/\sigma_{AL}$ sources are preferentially accepted, while for the 7-parameter solutions, lower $a_{0}/\sigma_{AL}$ are preferentially accepted.
Figure 16: Top four panels: Significance of the 9-parameter acceleration solution (left column) and parallax over error (right column), as a function of true orbital period vs. eccentricity (top row), and orbital period vs. photocenter orbit semimajor axis (bottom row), for fixed noise (i.e., magnitude and visibility periods). Bottom four panels: Same as the top four panels, but for the 7-parameter solutions. For both the 9-parameter and 7-parameter acceleration solutions, the significance and parallax over error are strong functions of the orbital period and the size of the photocenter orbit semimajor axis, but do not depend on the orbital eccentricity.
Figure 17: Probability of acceptance of the 9-parameter (left) and 7-parameter acceleration solutions (right), as a function of the true orbital period and the photocenter orbit semimajor axis, scaled by the per-observation uncertainty $\sigma_{AL}$ (which is a function of magnitude). The probability of 9-parameter solution acceptance is relative to all sources that have $\texttt{ruwe}>1.4$; the probability of 7-parameter solution acceptance is relative to all sources that did not have a 9-parameter solution accepted.
Appendix D Cross-match to external catalogs